OpenStudy (anonymous):

int \frac{ e^3x }{ (1+e^3x)^8 } *dx

OpenStudy (anonymous):

is it let u = 1 + e^3x ?

OpenStudy (anonymous):

\[\int\limits \frac{ e^3x }{ (1+e^3x)^8 } *dx \]